Week of 160530

WLCG Operations Call details

  • At CERN the meeting room is 513 R-068.

  • For remote participation we use the Vidyo system. Instructions can be found here.

General Information

  • The purpose of the meeting is:
    • to report significant operational issues (i.e. issues which can or did degrade experiment or site operations) which are ongoing or were resolved after the previous meeting;
    • to announce or schedule interventions at Tier-1 sites;
    • to inform about recent or upcoming changes in the experiment activities or systems having a visible impact on sites;
    • to provide important news about the middleware;
    • to communicate any other information considered interesting for WLCG operations.
  • The meeting should run from 15:00 until 15:20, exceptionally to 15:30.
  • The SCOD rota for the next few weeks is at ScodRota
  • General information about the WLCG Service can be accessed from the Operations Web
  • Whenever a particular topic needs to be discussed at the daily meeting requiring information from site or experiments, it is highly recommended to announce it by email to wlcg-operations@cernSPAMNOTNOSPAMPLEASE.ch to make sure that the relevant parties have the time to collect the required information or invite the right people at the meeting.

Tier-1 downtimes

Experiments may experience problems if two or more of their Tier-1 sites are inaccessible at the same time. Therefore Tier-1 sites should do their best to avoid scheduling a downtime classified as "outage" in a time slot overlapping with an "outage" downtime already declared by another Tier-1 site supporting the same VO(s). The following procedure is recommended:
  1. A Tier-1 should check the downtimes calendar to see if another Tier-1 has already an "outage" downtime in the desired time slot.
  2. If there is a conflict, another time slot should be chosen.
  3. In case stronger constraints cannot allow to choose another time slot, the Tier-1 will point out the existence of the conflict to the SCOD mailing list and at the next WLCG operations call, to discuss it with the representatives of the experiments involved and the other Tier-1.

As an additional precaution, the SCOD will check the downtimes calendar for Tier-1 "outage" downtime conflicts at least once during his/her shift, for the current and the following two weeks; in case a conflict is found, it will be discussed at the next operations call, or offline if at least one relevant experiment or site contact is absent.

Links to Tier-1 downtimes




  • local: Maria Alandes (Chair and minutes), Maarten Litmaath (ALICE), Andrea Manzi (MW Officer), Oliver Gutsche (CMS), Ben Jones (Computing), Jesús López (Storage), Kate (DB), Maria Dimou (GGUS, Security), Raja (LHCb)
  • remote: Eric Lancon (BNL), Francesco Noferini (CNAF), Dave Mason (FNAL), Rolf Rumler (IN2P3), Victor Zhiltsov (JINR), Xavier Mol (KIT), Onno Zweers (NL-T1), Pepe Flix (pic), Di Qing (TRIUMF), Vincenzo Spinoso (EGI)

Experiments round table:

  • CMS reports (raw view) -
    • data taking resumed Thursday afternoon, mediocre success
    • Discussion to postpone/cancel Machine Development 1 and Technical Stop 1 planned for Week 23 to maximize luminosity for ICHEP
      • MD1 block postponed to (at least) mid-July.
      • TS1 block: official decision tomorrow.
    • MC DigiReco for ICHEP continuing
    • Saturday, TWiki including shifter instructions were not reachable
      • alarm ticket GGUS:121817
      • Twiki NFS server was down. Kernel lockup due to ZFS module. Module was upgraded to latest version before hard rebooting.
      • Thanks for fixing on a Saturday!

Oli adds that there is a read only version of the twiki that is running in Wigner which could be used when the normal twiki is down. Maarten explains that there is in fact a failover mechanism that should switch to read only twiki but that didn't work this time. Maria adds that Pete Jones is in fact investigating why the failover switch didn't work this time.

  • ALICE -
    • mostly high activity

  • LHCb reports (raw view) -
    • Activity
      • Most user and simulation jobs on the grid
    • Site Issues
      • T0: Some problems transferring data from online to offline. Working on it.
      • T1: RAL : tape systems at risk, SARA : tape system downtime till tomorrow afternoon.

Maria asks what it means exactly that the RAL tape systems are at risk. Raja explains that this is related to RAL's report as explained below in the minutes and it refers to the HW problems RAL is experiencing where the tape mount fails when data is read.

Sites / Services round table:

  • ASGC: NA
  • BNL: NTR
  • GridPP: NA
  • IN2P3: NTR
  • KIT: NTR
    • Announced the fourth attempt to our tape maintenance downtime, where we will migrate the core management service onto new hardware, for 28th/29th of June.
  • NDGF: NA
  • NL-T1:
    • SARA's mass storage system currently in maintenance, until Wednesday.
    • SARA's datacenter move announcement on website: https://userinfo.surfsara.nl/movedc/grid-facilities
      • First 3 weeks of August: mass storage system moved
      • First 2 weeks of October: other grid systems are moved.
  • NRC-KI: NA
  • OSG: NA
  • PIC: NTR
  • RAL:
    • RAL is closed today (30th May). It is a holiday in the UK.
    • There have been problems with the control software for the tape library. This caused problems for tape mounts Wednesday/Thursday 25/26 May. There are a couple of hardware faults on the library. These do not affect services but an engineer visit is planned for Wednesday 1st June. An outage of the tape system is expected for a couple of hours on that day.

  • CERN computing services: NTR
  • CERN storage services:
    • ATLAS EOS will be updated if eventually there is a technical stop.
    • FTS: an issue affecting only CMS Phedex transfers has been discovered last week. ( FTS Optimizer not enabled on some links). The problem has been fixed and also a workaround has been applied to enable the optimizer for the missing links. The fixed release will be deployed Wed 1 Jun from 9:30 to 10:30 and the upgrade will be transparent to clients.
  • CERN databases: OS updates will be applied to all hosts running production databases next week (provided that the technical stop is confirmed). Detailed schedule is still being discussed with the experiments.
  • GGUS: (MariaD) Release this Wed 1st of June with ALARM tests as usual. NB! GGUS server security patches installation including kernel changes that may lead to some unavailability of the service JIRA:GGUS-1523.
  • Monitoring:
    • SAM BDII will be under risk due to migration of VMs 1st of June 9am-1pm, this could impact SAM/ETF job submissions
    • Draft reports for the May availability sent around
  • MW Officer:
    • edg-mkgridmap v4.0.2 does not work on CentOS/EL/SL 6.8
      • please update to v4.0.3 available in the WLCG rpm repository
      • EPEL6 will be updated with that version ASAP
      • it is backward-compatible with earlier OS releases
  • Networks:
    • GGUS:119820 ASGC connectivity. After numerous tests performed in collaboration with ASGC and ESNet (http://etf.cern.ch/perfsonar_asgc.txt) the root cause has been confirmed to be the local N7K router at ASGC. Once the perfSONARs were moved directly to the central router the measured network performance has improved by factor 10. Our recommendation is to re-wire all the existing data transfer nodes to bypass the local router as well as to tune the central router and data transfer nodes to improve their performance for long path transfers (200ms+).
    • GGUS:121687 RAL consistent loss - waiting for an upgrade of the router at RAL
  • Security: (MariaD) Updated several pages under http://wlcg.web.cern.ch/getting-started/certificates to reflect changes around the OSG PKI. Experiments, please inform MariaD via e-group project-lcg-certificates at cern dot ch (which appears on these web pages) if new sites join the grid and have not yet a national CA.


  • Vincenzo announces that the new version of edg-mkgridmap will be distributed in UMD in the next days.
  • Raja asks whether there is any clear timeline to fix the problem reported in GGUS:120586. Vincenzo explains there is no clear date yet but that he will follow up with CREAM developers.
Edit | Attach | Watch | Print version | History: r21 < r20 < r19 < r18 < r17 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r21 - 2016-06-01 - BorjaGarridoBear
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright &© 2008-2021 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
or Ideas, requests, problems regarding TWiki? use Discourse or Send feedback